### v1

# Benchmarking Large Language Models with Integer Sequence Generation Tasks

Abstract
This paper presents a novel benchmark where the large language model (LLM) must write code that computes integer sequences from the Online Encyclopedia of Integer Sequences (OEIS), a widely-used resource for mathematical sequences. The benchmark is designed to evaluate both the correctness of the generated code and its computational efficiency. Our benchmark reveals that the o1 series of models outperform other frontier models from OpenAI, Anthropic, Meta, and Google in accuracy and cheating rates across both easy and hard integer sequences. In order to ensure models do not exploit memorized sequence values, we introduce an automated cheating detection mechanism that flags the use of lookup tables and validated this automation against human cheating evaluations. This benchmark provides a meaningful challenge for current LLMs, offering insights into their mathematical reasoning and code writing capabilities, which can guide future research directions and model development in mathematical reasoning and code synthesis.
1 Introduction
Benchmarking plays a crucial role in the development and evaluation of large language models (LLMs), helping researchers and users gauge their abilities across various domains such as natural language understanding, knowledge retrieval, and mathematical reasoning. The progress that LLMs have made on challenging benchmarks is remarkable—matching even the performance of humans with a Ph.D. on advanced problems in the human’s domain of expertise. With the release of more powerful models like OpenAI’s o1 series, there is a need for benchmarks that can rigorously test more advanced abilities of these systems.
In this paper, we introduce a novel benchmark based on integer sequence generation tasks sourced from the Online Encyclopedia of Integer Sequences (OEIS) [1, 2]. OEIS provides a wealth of information on these sequences under a CC BY-SA 4.0 license, including sequence values that can be used for unit tests as well as sequence definitions and commentary that provide the LLM with information about the sequence. The difficulty of these tasks ranges from trivial (A000004 is the sequence of all zeros) to extremely difficult and interesting (A000001 is the number of groups of order n – “a fascinating function” for which Conway et al. [3] recently provided only an approximation of the series). These tasks are particularly challenging for LLMs, as they require the models not only to understand the sequences but also to implement efficient algorithms that can run quickly (both to test the algorithms and for the expediency of the benchmark). This makes integer sequence generation an excellent testbed, especially for models like o1-mini and o1-preview, which are optimized for tasks that demand reasoning skills such as mathematics and coding.
Our benchmark consists of a diverse set of 500 integer sequences labeled “easy” and “hard” in OEIS. We evaluate o1-preview [4], o1-mini [5], GPT-4o [6], GPT-4o-mini, and Claude 3.5 Sonnet [7], Llama 3.1 405b and 70b [8], and Gemini 1.5 [9] Pro and Flash on their ability to write code that generates these sequences. The codes are subject to a time limit that is allowed to vary (similar to a pass@k metric, where different values of k are used), analyzing their performance in terms of both accuracy and efficiency. While o1 models, with their reasoning capabilities, outperform the other models, it still struggles, especially with the hard sequences. Additionally, we introduce mechanisms for detecting and preventing the use of lookup tables to verify that models write legitimate code rather than relying on a memorized version of the sequence entries. See Figure 1 for an overview of our approach.
Our contributions are as follows: (1) We introduce a new benchmark for LLMs based on integer sequence generation, emphasizing mathematical and computational reasoning and efficiency. (2) We evaluate numerous frontier LLMs, demonstrating their strengths and limitations in handling these algorithmic tasks. (3) We provide a framework for detecting and mitigating the use of lookup tables in sequence generation tasks, bolstering the integrity of the evaluation process.
Figure 1: The process for curating easy and hard sequences, using an LLM to generate code that computes the sequence values, and evaluating the code is shown.
2 Related Work
Benchmarking has been essential in evaluating the capabilities of LLMs across various domains, particularly in mathematical reasoning and code generation. Existing benchmarks such as MATH [10], GSM8K [11], and HumanEval [12] assess models on complex problem-solving and programming tasks. While these benchmarks provide valuable insights, they often either cover broad problem areas or focus on specific aspects like functional correctness in code generation [12, 13]. Our benchmark distinguishes itself by concentrating on algorithmic reasoning through integer sequence generation, demanding both mathematical insight and efficient code implementation. This approach enables a deeper evaluation of LLMs’ capacity to generate accurate, efficient algorithms, filling a gap in existing benchmarks by challenging the latest, most advanced models in a meaningful way.
3 Benchmark Design
In designing the benchmark, our goal was to create a robust and rigorous evaluation framework that challenges frontier LLMs. The benchmark is centered around writing code that computes elements of integer sequences, using sequences sourced from the OEIS [1]. The design incorporates various levels of difficulty and enforces performance constraints to measure both the accuracy and efficiency of the models.
3.1 Dataset Selection
The dataset for the benchmark is derived from OEIS, an extensive database of integer sequences contributed by mathematicians worldwide. We selected 250 easy and 250 hard sequences based on OEIS labels. The set of sequences is defined as $\mathcal{S}=\mathcal{S}_{\text{easy}}\cup\mathcal{S}_{\text{hard}}$, where $\mathcal{S}_{\text{easy}}$ are the first 250 sequences labeled as easy, and $\mathcal{S}_{\text{hard}}$ are the first 250 sequences labeled as hard in OEIS.
This selection provides a broad spectrum of sequence generation problems, ranging from basic arithmetic operations to complex mathematical computations.
3.2 Problem Definition
For each sequence $s\in\mathcal{S}$, an LLM $M$ is tasked with generating Python code $C_{s}$ that computes the first $N$ terms of the sequence $s$, where $N$ is a fixed positive integer (e.g., $N=10$). Each integer sequence is a function:
$s:\{i_{0}+j\}_{j=0}^{\infty}\rightarrow\mathbb{Z},$
where $i_{0}$ is an offset indicating where the sequence starts. The code $C_{s}$ should define a function
$f_{s}:\{i_{0}+j\}_{j=0}^{\infty}\rightarrow\mathbb{Z}$
such that $f_{s}(n)=s(n)$ for all $n\geq i_{0}$.
The following constraints are imposed on the generated code: (1) the code $C_{s}$ must not contain a lookup table of the sequence terms, (2), the execution time $t_{s}$ of $C_{s}$ must satisfy $t_{s}\leq T$ where $T$ is a predefined time limit (we use $T\in\{0.5,4\}$ seconds), and (3) the code must be valid Python code executable in a standard environment without external library dependencies.
3.3 Evaluation Metrics
To provide a comprehensive evaluation of the models, we measure their performance using three factors: accuracy, efficiency, and avoiding lookup tables.
For each sequence $s$, we define the accuracy $A_{s}(n)$ as:
$A_{s}(n)=\begin{cases}0&f_{s}(n)\neq s(n)\\ 0&t_{s}>T\\ 0&\mathrm{cheating}\\ 1&\mathrm{otherwise}\end{cases}$ (1)
We report the average accuracy over all sequence values in $\mathcal{S}_{\text{easy}}$ and $\mathcal{S}_{\text{hard}}$. Note that $f_{s}(n)$ must execute within the time limit $T$ or the delta fun
3.4 Cheating Detection Mechanism
Another core aspect of the benchmark is ensuring that models produce algorithms rather than lookup tables of sequence values. To enforce this, we use GPT-4o’s structured output function (with temperature 0 to maximize reproducibility) to check the code output by the model and flags cases where lookup tables are employed. Any model that is found to be cheating by using a lookup table receives a score of zero for that task, regardless of the accuracy of the output. This cheating detection mechanism’s effectiveness was validated by comparing it with a human evaluation (one of the authors, who was not provided with the GPT-4o cheating evaluations beforehand). The human was provided the same instructions as GPT-4o, which are provided in the supplementary information. The automated approach achieved 86% agreement overall and 94% agreement on hard sequences, where cheating is more prevalent. The main mode of disagreement between the human evaluator and GPT-4o is that GPT-4o flags more codes as cheating.
4 Experiments and Results
We evaluate nine state-of-the-art LLMs on our integer sequence benchmark using their default settings (temperature, etc.). Table 1 summarizes the models’ performance on the easy ($\mathcal{S}_{\text{easy}}$) and hard ($\mathcal{S}_{\text{hard}}$) sequence sets. Figure 2 shows more details about the distribution of scores. There are only small differences when the 0.5s and 4s time limits are used, so we focus our discussion on the 4s case. The o1 models give the best performance for both easy and hard sequences, scoring 63% or higher on the easy sequences and 18% or higher on the hard sequences. The o1-mini model performs slightly better than o1-preview on the easy sequences and vice-versa for the hard sequences. The o1 models have a clear performance gap compared to the non-reasoning models, especially on the hard sequences where they almost double the performance of strong non-reasoning models like Claude 3.5 Sonnet and GPT-4o. The best non-reasoning model is Claude 3.5 Sonnet, which scores 57% on easy sequences and 11% on the hard sequences. Multiple frontier models score below 50% on the easy sequences (Llama 405b and Gemini 1.5 Pro) and below 10% on the hard sequences (GPT-4o, Llama 405b and Gemini 1.5 Pro).
Model Time SequenceEasy SequenceHard
Avg. Score Cheating Avg. Score Cheating
o1-mini 0.5 65.3 2.0% 17.3 15.2%
4 65.4 2.0% 18.1 15.2%
o1-preview 0.5 63.2 2.8% 18.0 23.2%
4 63.2 2.8% 18.7 23.2%
claude-3.5-sonnet 0.5 56.9 2.0% 10.8 42.0%
4 57.0 2.0% 11.1 42.0%
gpt-4o 0.5 54.5 4.0% 8.7 37.2%
4 54.6 4.0% 8.9 37.2%
gpt-4o-mini 0.5 51.6 6.0% 8.5 32.0%
4 52.0 6.0% 8.7 32.0%
llama-405b 0.5 47.4 4.4% 9.9 32.0%
4 47.6 4.4% 10.1 32.0%
gemini-1.5-pro 0.5 49.4 20.8% 4.7 73.2%
4 49.4 20.8% 4.9 73.2%
llama-70b 0.5 44.7 2.4% 8.6 22.0%
4 44.7 2.4% 8.7 22.0%
gemini-1.5-flash 0.5 45.9 4.0% 6.5 36.0%
4 46.1 4.0% 6.6 36.0%
Table 1: The performance of frontier models on $\mathcal{S}_{\text{easy}}$ and $\mathcal{S}_{\text{hard}}$ is shown. The average score includes zeros for instances where cheating was detected.
Figure 2: The distribution of scores is shown, where the sequences are sorted for each sequence set ($\mathcal{S}_{\text{easy}}$ and $\mathcal{S}_{\text{hard}}$) and model pair according to the LLM’s accuracy computing the sequence.
Our cheating detection mechanism used GPT-4o’s structured outputs to identify instances where models used lookup tables to compute sequences. The results show that o1-mini has the lowest cheating rate, with $2\%$ in $\mathcal{S}_{\text{easy}}$ and $15.2\%$ in $\mathcal{S}_{\text{hard}}$. Gemini 1.5 Pro is heavily penalized for using lookup tables in $\mathcal{S}_{\text{hard}}$ and scores below 5%. The Gemini API frequently refused to return results for the prompts in this benchmark because a filter determined that it was producing a copyrighted result. This may have contributed to frequent use of lookup tables and low score.
5 Discussion
The superior performance of o1-mini highlights the effectiveness of specialization in LLMs for mathematical reasoning and coding tasks. o1-mini’s higher accuracy and lower cheating rates demonstrate that models optimized for STEM reasoning can significantly outperform general-purpose models on algorithmic tasks. The low average scores on $\mathcal{S}_{\text{hard}}$ across all models indicate that current LLMs struggle with generating complex algorithms, emphasizing the need for further advancements in this area. The higher cheating rates in GPT-4o suggest that models may default to memorization when faced with difficult tasks. Improving models’ ability to derive algorithms rather than recalling answers is crucial.
Using integer sequence generation from OEIS has proven to be an effective benchmark for evaluating computational reasoning in LLMs. The richness of the OEIS dataset, with its diverse range of sequences, provides a challenge for models across varying levels of difficulty. The sequences test both basic and advanced mathematical concepts, making them ideal for evaluating LLMs’ mathematical reasoning and code writing. Additionally, the benchmark was cost-effective to evaluate, requiring around \$200 in API credits (mostly for the o1 models), making it accessible for broad use.
Implementing cheating detection mechanisms was instrumental in developing an effective benchmark, given the frequent occurrence of lookup tables in $\mathcal{S}_{\text{hard}}$. By identifying when models used lookup tables, we ensured that the benchmark tested their ability to generate solutions algorithmically. This mechanism played a key role in maintaining the integrity of the evaluation.
There are several avenues for future research. Integrating tool use, such as web access combined with retrieval-augmented generation (RAG), could enable models to access additional resources during problem-solving. Allowing LLMs to retrieve and utilize external information, like the extensive references and comments available in OEIS entries, may improve their ability to generate algorithms for complex sequences. In this study, such reference information was not provided to the models. Future benchmarks could incorporate these resources to assess models’ abilities to leverage external knowledge effectively. Including newly added OEIS sequences in the benchmark can maintain its relevance and challenge as models advance. The OEIS database is continuously updated with new sequences less likely to have been encountered during model training. Testing models on these sequences would evaluate their capability to generalize to novel problems. However, given the current models’ low performance on existing sequences—particularly the hard ones—immediate expansion may not be essential. Enhancing model performance on the current dataset should be a priority before introducing additional sequences.
6 Conclusion
We introduced a rigorous benchmark to evaluate large language models on generating code for integer sequences from the OEIS, focusing on mathematical reasoning and computational efficiency. Our evaluation demonstrated that reasoning models like o1-mini o1-preview outperform general-purpose models such as GPT-4o, Claude 3.5 Sonnet, and Gemini 1.5 Pro in tasks requiring mathematical insight and algorithmic coding skills. Specifically, the o1 models achieved higher accuracy and lower cheating rates on both the easy ($\mathcal{S}_{\text{easy}}$) and hard ($\mathcal{S}_{\text{hard}}$) sequence sets. Despite these strengths, all models showed low performance on the hard sequences, underscoring the challenges LLMs face in generating complex algorithms within practical time constraints. The frequent reliance on memorization strategies, like using lookup tables, highlights the need for developing models capable of genuine algorithmic reasoning. Our benchmark effectively assessed the computational reasoning abilities of LLMs, with the OEIS dataset providing a robust and diverse evaluation framework. The implemented cheating detection mechanism was crucial in ensuring adherence to algorithmic constraints and maintaining the integrity of the assessment.
7 Limitations
Our benchmark possesses several limitations that warrant consideration. First, relying exclusively on the OEIS as the source of integer sequences may introduce biases due to the specific types and distributions of sequences included, as well as their subjective labeling as “easy” or “hard.” Another potential issue is that some OEIS sequences have associated code snippets that are publicly available. While this could assist LLMs in generating the correct sequence, many of the selected problems are difficult enough to require novel mathematical insights or optimized coding techniques. The codes that are available in the OEIS database are most often in languages like Mathematica, Maple, or Magma and tend to rely on advanced functionality of these pieces of software that can turn complex sequences into a few lines of code. To mitigate this, we require the models to generate code using the Python standard library, which does not have equivalent functionality. In practice, current models face considerable challenges in computing the sequences efficiently, especially for the $\mathcal{S}_{\text{hard}}$ problems. The OEIS is also actively maintained with 30-60 new sequences being added on a daily basis [14]. So, the benchmark could be continuously updated to mitigate the effect of OEIS sequence information being included in the training of LLMs.
Second, although our cheating detection mechanism effectively identifies the use of lookup tables, it is not infallible. With an $86\%$ agreement rate with human evaluators, some instances of cheating may go undetected or be falsely flagged, potentially impacting the accuracy of the evaluation. Instances where the human flagged the code as not cheating and GPT-4o flagged it as cheating where by far the most common discrepancy (see Table 2). Actually, instances of cheating that were flagged by the human but not by GPT-4o did not occur in our evaluation (though certainly they could occur). Discrepancies between automated and human assessments indicate a potential for further refinement of the detection method. This could be done by using other LLMs (e.g., o1-mini might do better than GPT-4o) or using a human to do the full cheating analysis. Using a human, of course, has the drawback of making the benchmark much more expensive to evaluate. A possible middle ground would be to have a human evaluate a subset of codes for cheating, then use few-shot in-context learning to align the evaluations of the LLM with the human evaluations. Of course, there is fundamentally some subjectivity in the determination of whether or not a code uses a lookup table.
Cheating? GPT-4o: Yes GPT-4o: No
Human: Yes 54 0
Human: No 26 90
Table 2: Confusion matrix for our evaluation of the GPT-4o based cheating detection, as compared with an expert human evaluator.
Third, restricting code generation to Python confines the evaluation to a single programming language. This limitation may not fully capture a model’s versatility or efficiency in other languages that could be more suitable for certain sequences. Models might perform differently if allowed to utilize languages better aligned with the computational demands of specific tasks. Python’s has a relatively slow computational speed. At the same time, many sequences have a seemingly high cost to compute some of the sequence values. For example, A000791 [15] is a sequence in $\mathcal{S}_{\mathrm{hard}}$ of Ramsey numbers – notoriously difficult to compute – and one of the comments notes that the tenth element in the sequence being 42 was ruled out “with a massive computer search.” This combination of a slow language and expensive computations suggests that a different program language might provide the models a better chance of success.
Fourth, the imposed time constraints, while essential for assessing efficiency, may disadvantage models that implement correct but computationally intensive algorithms, especially for sequences inherently requiring significant resources. This could unfairly penalize models due to factors beyond their control, such as hardware limitations or the intrinsic complexity of the problem.
Last but not least, the OEIS is a rich, human-curated resource that does not accept contributions from artificial intelligence models. Therefore, it would be unethical for LLM-generated codes to be upstreamed into the OEIS. Avoiding this also serves the goals of this benchmark by not entering a loop where LLMs are trained on OEIS data that was generated by LLMs.
Supplementary Information
Here, we provide the prompts we used for code generation and lookup table detection using Python’s string interpolation syntax. The actual prompts would contain the interpolated information.
Code Generation Prompt
[…]
Lookup Table Detection Prompt
[…]